# Strategic Investment to Mitigate Transition Risks

### v1

@@ 2.3.3 Sensitivity Analysis
Finally, the analysis of $B$ shows that as carbon prices increase from 0.50 to 1.50, the optimal alpha rises from 0.24 to 0.49, suggesting that the firm needs to allocate more resources to low-carbon production as carbon prices rise. However, this comes at a cost, as the maximum objective value declines from 63.89 to 39.91, showing that higher carbon prices can lead to a reduced level of profitability.
@@
The change rate of alpha is consistently negative, starting at $-35.75\%$ for $B=0.50$ and improving slightly to $-6.74\%$ at $B=1.10$, indicating that while increased carbon pricing necessitates the firm to invest more, its negative impact on profitability is relatively moderate. The change rate of the profit peaks at $26.67\%$ for $B=0.50$ and drops to $-20.87\%$ at $B=1.50$. This suggests that while carbon pricing does affect firm’s profitability, the sensitivity is not as severe for the firm compared to other variables. Companies can mitigate the impact of higher carbon prices through improved efficiency and strategic investments in cleaner technologies, ensuring that their profitability does not drastically diminish as carbon prices rise.
@@
In comparing the variables in the framework, it is clear that the original productivity coefficient $(\beta)$ is by far the most sensitive. Changes in the value of $\beta$ lead to substantial shifts in the optimal value of alpha, maximum objective value, and the change rate of the profit. The strong relationship between productivity and profitability suggests that firms should prioritize improving productivity to maximize financial outcomes.
@@
Conversely, the carbon price per unit $(B)$ is the least sensitive variable in the framework. While changes in $B$ do affect both the optimal alpha and the profitability, the overall impact for the firm is moderate compared to the other factors. Firms can manage the effects of carbon pricing through strategic investments in cleaner technologies, making it a more controllable factor in its financial performance.
@@
2.3.4 Expanding the One-Period Model to a Multi-Period Model
In our initial analysis, we have focused on a static one-period model to evaluate the impact of transition investment on enterprise profitability. This framework provided basic but useful insights into the immediate effects of various variables on a company’s profit. However, real-world scenarios are dynamic in nature and involve changes over multiple periods. Therefore, as a next logical step, we are expanding our previous static framework to a multi-period setting to capture the evolving nature of the variables in the framework and their long-term impacts.
To achieve this, we revise the profit function to account for time dependency, reflecting the changes in key variables in the framework over multiple different periods. The time-dependent profit function is defined as follows:
@@ 2.3.4 Expanding the One-Period Model to a Multi-Period Model
This formulation captures the dynamic nature of transition investments, where $p$ is the selling price per unit (assume $p$ is a constant), $c_{t}$ is the production cost per unit at time $t,\alpha$ is the transition investment ratio, $k_{t}$ is the enterprise low-carbon production efficiency coefficient at time $t,B_{t}$ is the carbon price at time $t$, and $\beta_{t}$ is the original productivity coefficient at time $t$. By incorporating these time-dependent variables, we can more accurately model the impact of transition investments over multiple periods.
The resulting multi-period objective function can then be expressed as follows:
@@
where $g$ is the discount factor. This objective function accounts for firm’s profits in the current period $\left(\pi_{t}\right)$ and its discounted future profits ($\pi_{t+1},\pi_{t+2}$, and so on), providing a more complete view of the impact of transition investments for the firm. The inclusion of future profits, discounted to the present time by the factor $g$, acknowledges the time value of money, opportunity costs, and inherent risks faced by the company. The discounting factor also acknowledges the uncertainties associated with long-term projections, ensuring a realistic assessment of the enterprise’s financial trajectory.
By using this multi-period framework, we aim at finding the optimal value of $\alpha$ that maximizes the overall profit of the firm across multiple periods. The multi-period model is particularly useful for understanding the long-term implications of transition investments for the firm. Transitional investments typically incur immediate costs but generate substantial future benefits for the firm. For example, investments in low-carbon technologies might initially increase production costs for the firm due to the capital expenditure required for new equipment or processes. However, over time, these investments can lead to substantial efficiency gains, cost reductions, and improved regulatory compliance, which enhance future profits of the company.
Short-Term vs. Long-Term Observation:
Our goal is to find the optimal value of $\alpha$ to achieve the maximum objective function value for the firm over different observation periods. For simplicity, we will compare a shorter-term observation (of 3 years) with a longer-term observation (of 6 years).
  • Short-Term Observation (3 years): This period allows us to understand the immediate and near-future impacts of transition investments for the firm. It is suitable for enterprises looking for quick wins and immediate adjustments in their strategies.
  • Long-Term Observation (6 years): This extended period captures the long-term benefits and potential drawbacks of transition investments faced by the company. It helps the enterprises’ plan strategically for sustained growth and ensures compliance with evolving regulations and market conditions.
By comparing these two observation periods, we can evaluate how the optimal value of $\alpha$ varies based on the length of the planning horizon for the company. This comparison will provide important insights into the trade-offs between short-term gains and long-term sustainability, guiding enterprises in making informed decisions about their transitional investments. The numerical exercise will be presented in the next subsection to illustrate these findings.
@@ 2.4 Decarbonization Scenario Analysis
2.4.1 Introduction to Decarbonization Scenarios
To illustrate the practical application of our multi-period framework, we examine four decarbonization scenarios by means of a numerical method: Immediate, Quick, Slow, and No Decarbonization. Decarbonization scenarios are strategic frameworks used to explore and analyze the pathways by which industries, economies, and societies can make a transition from high-carbon to low-carbon or carbon-neutral operations. These scenarios are designed to address the urgent need to mitigate climate change by reducing GHG emissions, particularly carbon dioxide (CO2), which is viewed as a major contributor to human-induced global warming. Each decarbonization scenario represents a different approach, pace, and scale of implementing low-carbon technologies, policies, and practices.
These scenarios draw inspiration from the pathways outlined in IPCC’s Representative Concentration Pathways (RCPs) and Shared Socioeconomic Pathways (SSPs) (IPCC (2018, 2022)), which provide a framework for modeling emissions trajectories under varying socio-economic, policy, and technological conditions. For example, the Immediate Decarbonization scenario aligns with the stringent measures seen in RCP1.9/SSP1-1.9, where immediate, large-scale investments aim to limit global temperature rise to 1.5°C (IPCC (2018)). Conversely, the No Decarbonization scenario parallels RCP8.5/SSP5-8.5, representing a ”business-as-usual” trajectory with minimal climate action and escalating carbon emissions (IPCC (2023)).
@@ 2.4.1 Introduction to Decarbonization Scenarios
In this scenario, significant and immediate investments are made in low-carbon technologies and practices. The goal is to achieve rapid reductions in carbon emissions as quickly as possible. This approach mirrors the aggressive pathways described in RCP1.9/SSP1-1.9, where rapid decarbonization is pursued through stringent regulatory measures, substantial financial incentives, and rapid deployment of renewable energy and energy efficiency technologies (IPCC (2018)). Immediate decarbonization is typically driven by strong political will and societal urgency to address climate change impacts swiftly. While effective in achieving fast emissions reductions, it may entail higher short-term costs and risks of over-investment.
Quick Decarbonization:
@@
Quick Decarbonization involves a fast-paced yet balanced transition. Investments in low-carbon technologies are substantial but spread out over a few years rather than all at once. This scenario resembles the pathways described in RCP2.6/SSP1-2.6, where emissions reductions are significant but distributed over time to balance economic feasibility and technical constraints (IPCC (2022); Agency (2022)). Quick decarbonization includes strategies such as accelerated deployment of renewable energy, electrification of transportation, and enhanced energy efficiency measures. It offers a cost-effective and pragmatic approach to achieving climate goals while reducing the economic and logistical burden of an immediate overhaul.
Slow Decarbonization:
@@
The Slow Decarbonization scenario adopts a more gradual approach to transitioning to a low-carbon economy. Investments and regulatory changes are implemented over a longer period of time, allowing for a smoother transition with less immediate economic disruption. This approach corresponds to RCP4.5/SSP2-4.5, a “middle-of-the-road” pathway where moderate policies and socio-economic factors lead to incremental emissions reductions over time (IPCC (2023)). While less disruptive, this scenario risks delaying critical action, potentially increasing long-term mitigation costs and reliance on future technological breakthroughs.
No Decarbonization:
@@
In the No Decarbonization scenario, no substantial efforts are made to reduce carbon emissions. This approach parallels the ”business-as-usual” trajectory represented by RCP8.5/SSP5-8.5, where industries and societies continue with their current high-carbon practices (IPCC (2018)). Continued reliance on fossil fuels results in escalating carbon emissions, environmental degradation, and increased climate risks (Agency (2022); IPCC (2023)). This scenario is widely regarded as unsustainable, as it amplifies the adverse economic, environmental, and social impacts of climate change over time.
2.4.2 Variable Changes and Rationale for Each Decarbonization Scenario
@@
To effectively analyze the aforementioned impacts of different decarbonization strategies, it is essential to understand how key variables in the framework evolve under each scenario and the rationale behind these changes. This part delves into the specific modifications of variables such as the enterprise low-carbon production efficiency coefficient $\left(k_{t}\right)$, production cost per unit ($c_{t}$), original productivity coefficient ($\left.\beta_{t}\right)$, and carbon price $\left(B_{t}\right)$ for each of the decarbonization scenarios: Immediate, Quick, Slow, and No Decarbonization. By doing so, we can better comprehend the short-term and long-term effects of the transition on firm’s profitability and identify the optimal transition investment ratio $(\alpha)$ for the firm in each case.
Immediate Decarbonization:
@@
In the Immediate Decarbonization scenario, enterprises make substantial investments in low-carbon technologies immediately. This aggressive approach aims at achieving rapid decarbonization and substantial reductions in carbon emissions in the shortest possible time. As a result, the enterprise’s low-carbon production efficiency coefficient $\left(k_{t}\right)$ starts at a very high level due to a substantial initial level of investment, reflecting major improvements in production efficiency of the firm. Over time, as the immediate benefits are realized, $k_{t}$ decreases gradually. The production cost per unit $\left(c_{t}\right)$ of the firm decreases quickly as economies of scale and technological efficiencies are realized from the new low-carbon technologies. The original productivity coefficient $\left(\beta_{t}\right)$ increases initially due to enhanced production capabilities of the firm but stabilizes over time as the effects of initial investments are more fully integrated. The carbon price ($B_{t}$) remains at a very high level due to stricter regulations and market adjustments that favor low-carbon outputs.
Quick Decarbonization:
@@
The Quick Decarbonization scenario involves rapid but slightly less aggressive investments compared to the Immediate scenario. Enterprises aim at achieving substantial decarbonization within a relatively short time frame but spread their investments over a few periods of time to balance costs and benefits. Consequently, the enterprise’s low-carbon production efficiency coefficient $\left(k_{t}\right)$ increases quickly but not as sharply as in the case of the Immediate scenario. The efficiency gains are substantial for the firm but spread over a slightly longer period of time. The production cost per unit $\left(c_{t}\right)$ decreases rapidly but not as dramatically, reflecting more moderate cost reductions. The original productivity coefficient $\left(\beta_{t}\right)$ shows a steady increase due to technological improvements and better resource allocation by the firm. The carbon price $\left(B_{t}\right)$ keeps at a relatively high level, driven by increasing regulatory pressures and market responses.
Slow Decarbonization:
@@
In the Slow Decarbonization scenario, enterprises adopt a gradual approach to transitioning to low-carbon technologies. This approach spreads investments over a longer period of time, allowing enterprises to manage costs and mitigate risks while steadily reducing carbon emissions. As such, the enterprise’s low-carbon production efficiency coefficient $\left(k_{t}\right)$ increases slowly, reflecting incremental improvements in production efficiency for the firm as investments are made gradually. The production cost per unit $\left(c_{t}\right)$ decreases steadily but slowly, indicating progressive cost reductions over time. The original productivity coefficient $\left(\beta_{t}\right)$ shows a gradual increase, correlating with incremental enhancements in productivity as new technologies are adopted by the firm over time. The carbon price $\left(B_{t}\right)$ is set as a relatively low level, aligning with gradual regulatory changes and market adjustments.
No Decarbonization:
@@
The No Decarbonization scenario assumes that enterprises do not make any major investments in low-carbon technologies. Instead, they continue with their current practices, resulting in minimal changes to carbon emissions and production efficiency. Consequently, there is no need to find the optimal transitional investment ratio under this scenario, as the transitional investment ratio, $\alpha$, should always be zero.
2.4.3 Numerical Study
@@
By analyzing these different scenarios, we can observe how different levels of investment and urgency in transitioning to low-carbon technologies affect key variables of the framework over time. This analysis helps us understand the trade-offs between immediate costs and long-term benefits, guiding enterprises in making informed decisions about their decarbonization strategies. The goal is to determine the optimal value of $\alpha$ that maximizes the overall profit across multiple periods, ensuring both short-term (3-year observation period) profitability and long-term (6-year observation period) sustainability. We will not consider the scenario of ‘No Decarbonization’ in this part, as there is no need to determine the optimal transitional investment ratio $\alpha$ under this scenario.
@@
In the Quick Transition scenario, enterprises adopt low-carbon technologies at a moderately fast pace. This approach aims at achieving a rapid improvement in production efficiency, though at a slightly slower rate than in the Immediate scenario. As a result, the initial efficiency gains are significant but not as drastic, allowing the enterprise to avoid potential issues related to over-investment. Given the initially low value of $k$ (starting at 1.5), there is an urgent need to enhance production efficiency quickly. To address this, the enterprise makes a transition investment before the second year, resulting in a sharp increase in $k$ to 5 by the second year. This value then gradually declines each subsequent year, reflecting the natural efficiency losses due to wear and tear. Production costs per unit begin at a relatively high level (2.4), creating a strong need for transition investment to reduce costs. With the Quick Decarbonization approach, $c$ drops significantly to 1.6 in the second year. However, it then rises slightly in the following years, likely due to normal depreciation and the ongoing costs associated with maintaining the new technology. Similarly, the original productivity coefficient $\beta$ starts at a lower level (0.6), emphasizing the necessity for improvement. Due to the transition investment, $\beta$ reaches a peak of 0.9 in the second year, reflecting increased productivity. Over time, it gradually decreases, aligning with the adaptation phase and the natural reduction in productivity gains. Carbon pricing $B$ is maintained at a high but stable level (3) in the Quick Decarbonization scenario. This provides a continuous incentive for adopting low-carbon technologies, without imposing excessive financial burdens on the firm.

### v2

@@ A Total Asset 100
Carbon Price Per Unit $(B)$. Finally, the analysis of $B$ shows that as carbon prices increase from 0.50 to 1.50, the optimal $\alpha$ rises from 0.24 to 0.49, suggesting that the firm needs to allocate more resources to low-carbon production as carbon prices rise. However, this comes at a cost, as the maximum objective value declines from 63.89 to 39.91, showing that higher carbon prices can lead to a reduced level of profitability.
@@
The change rate of $\alpha$ is consistently negative, starting at $35.75\%$ for $B=0.50$ and improving slightly to $6.74\%$ at $B=1.10$, indicating that while increased carbon pricing necessitates the firm to invest more, its negative impact on profitability is relatively moderate. The change rate of the profit peaks at $26.67\%$ for $B=0.50$ and drops to $20.87\%$ at $B=1.50$. This suggests that while carbon pricing does affect firm’s profitability, the sensitivity is not as severe for the firm compared to other variables. Companies can mitigate the impact of higher carbon prices through improved efficiency and strategic investments in cleaner technologies, ensuring that their profitability does not drastically diminish as carbon prices rise.
@@
In comparing the variables in the model, it is clear that the original productivity coefficient $(\beta)$ is by far the most sensitive. Changes in the value of $\beta$ lead to substantial shifts in the optimal value of $\alpha$, the maximum objective value, and the change rate of the profit. The strong relationship between productivity and profitability suggests that firms should prioritize improving productivity to maximize financial outcomes.
@@
Conversely, the carbon price per unit $(B)$ is the least sensitive variable in the framework. While changes in $B$ do affect both the optimal $\alpha$ and the profitability, the overall impact for the firm is moderate compared to the other factors. Firms can manage the effects of carbon pricing through strategic investments in cleaner technologies, making it a more controllable factor in its financial performance.
@@
Initially, we have focused on a one-period model to evaluate the impact of transition investment on enterprise profitability. This static model provides basic but useful insights into the immediate effects of various variables on a company’s profit. However, real-world scenarios are dynamic in nature and evolve over time. Therefore, as a next logical step, we extend this one-period model to a multi-period setting to capture the evolving nature of the variables and some of the parameters in the model and their long-term impacts.
To achieve this goal, we revise the profit function to account for time dependency, reflecting the changes in key variables in the model over multiple periods. The time-varying profit function is defined as follows:
@@
This formulation captures the dynamic nature of transition investments, where $p$ is the selling price per unit (assume that $p$ is a constant), $c_{t}$ is the production cost per unit at time $t,\alpha$ is the transition investment ratio, $k_{t}$ is the enterprise low-carbon production efficiency coefficient at time $t,B_{t}$ is the carbon price at time $t$, and $\beta_{t}$ is the original productivity coefficient at time $t$. By incorporating these time-varying variables, we can more accurately model the impact of transition investments over multiple periods.
The resulting multi-period objective function can then be expressed as:
@@
where $g$ is a discount factor. This objective function accounts for firm’s profits in the current period $\left(\pi_{t}\right)$ and its discounted future profits ($\pi_{t+1},\pi_{t+2}$, and so on), providing a more complete view of the impact of transition investments for the firm. The inclusion of future profits, discounted to the present time by the factor $g$, acknowledges the time value of money, opportunity costs, and inherent risks faced by the company. The discounting factor also acknowledges the uncertainties associated with long-term projections, ensuring a more realistic assessment of the enterprise’s financial trajectory.
By using this multi-period framework, we wish to determine the optimal value of $\alpha$ that maximizes the overall profit of the firm across multiple periods. The multi-period model is particularly useful for understanding the long-term implications of transition investments for the firm. Transitional investments typically incur immediate costs but generate substantial future benefits for the firm. For example, investments in low-carbon technologies might initially increase production costs for the firm due to the capital expenditure required for new equipment or processes. However, over time, these investments can lead to substantial efficiency gains, cost reductions, and improved regulatory compliance, which enhance future profits of the company.
Short-Term vs. Long-Term Observation Periods. Our goal is to find the optimal value of $\alpha$ to achieve the maximum objective function value for the firm over different observation periods. For simplicity, we will compare a shorter-term period (of 3 years) with a longer-term period (of 6 years). The Short-Term Period allows us to understand the immediate and near-future impacts of transition investments for the firm. It is suitable for enterprises looking for quick wins and immediate adjustments in their strategies. The Long-Term Period captures the long-term benefits and potential drawbacks of transition investments faced by the company. It helps the enterprises’ plan strategically for sustained growth and ensures compliance with evolving regulations and market conditions over time.
By comparing these two periods, we can evaluate how the optimal value of $\alpha$ varies based on the length of the planning horizon for the company. This comparison will provide important insights into the trade-offs between short-term gains and long-term sustainability, guiding enterprises in making informed decisions about their transitional investments. The numerical exercise will be presented in the next subsection to illustrate these findings.
@@ 2.4 Decarbonization Scenario Analysis
To illustrate the practical application of our multi-period model, we examine four decarbonization scenarios by means of a numerical method: Immediate, Quick, Slow, and No Decarbonization. Decarbonization scenarios are strategic frameworks used to explore and analyze the pathways by which industries, economies, and societies can make a transition from high-carbon to low-carbon or carbon-neutral operations. These scenarios are designed to address the urgent need to mitigate climate change by reducing GHG emissions, particularly carbon dioxide (CO2), which is viewed as a major contributor to human-induced global warming. Each decarbonization scenario represents a different approach, pace, and scale of implementing low-carbon technologies, policies, and practices.
These scenarios draw inspiration from the pathways outlined in IPCC’s Representative Concentration Pathways (RCPs) and Shared Socioeconomic Pathways (SSPs) (IPCC (2018, 2022)), which provide a framework for modeling emissions trajectories under varying socio-economic, policy, and technological conditions. For example, the Immediate Decarbonization scenario aligns with the stringent measures seen in RCP1.9/SSP1-1.9, where immediate, large-scale investments aim at limiting a global temperature rise to 1.5°C (IPCC (2018)). Conversely, the No Decarbonization scenario parallels RCP8.5/SSP5-8.5, representing a ”business-as-usual” trajectory with minimal climate action and escalating carbon emissions (IPCC (2023)).
@@
Immediate Decarbonization. In this scenario, significant and immediate investments are made in low-carbon technologies and practices. The goal is to achieve rapid reductions in carbon emissions as quickly as possible. This approach mirrors the aggressive pathways described in RCP1.9/SSP1-1.9, where rapid decarbonization is pursued through stringent regulatory measures, substantial financial incentives, and rapid deployment of renewable energy and energy efficiency technologies (IPCC (2018)). Immediate decarbonization is typically driven by strong political will and societal urgency to address climate change impacts swiftly. While effective in achieving fast emissions reductions, it may entail higher short-term costs and risks of over-investment.
@@
Quick Decarbonization. Quick Decarbonization involves a fast-paced yet balanced transition. Investments in low-carbon technologies are substantial but spread out over a few years rather than all at once. This scenario resembles the pathways described in RCP2.6/SSP1-2.6, where emissions reductions are significant but distributed over time to balance economic feasibility and technical constraints (IPCC (2022); Agency (2022)). Quick decarbonization includes strategies such as accelerated deployment of renewable energy, electrification of transportation, and enhanced energy efficiency measures. It offers a cost-effective and pragmatic approach to achieving climate goals while reducing the economic and logistical burden of an immediate overhaul.
@@
Slow Decarbonization. The Slow Decarbonization scenario adopts a more gradual approach to transitioning to a low-carbon economy. Investments and regulatory changes are implemented over a longer period of time, allowing for a smoother transition with less immediate economic disruption. This approach corresponds to RCP4.5/SSP2-4.5, a “middle-of-the-road” pathway where moderate policies and socio-economic factors lead to incremental emissions reductions over time (IPCC (2023)). While less disruptive, this scenario risks delaying critical action, potentially increasing long-term mitigation costs and reliance on future technological breakthroughs.
@@
No Decarbonization. In the No Decarbonization scenario, no substantial efforts are made to reduce carbon emissions. This approach parallels the ”business-as-usual” trajectory represented by RCP8.5/SSP5-8.5, where industries and societies continue with their current high-carbon practices (IPCC (2018)). Continued reliance on fossil fuels results in escalating carbon emissions, environmental degradation, and increased climate risks (Agency (2022); IPCC (2023)). This scenario is widely regarded as unsustainable, as it amplifies the adverse economic, environmental, and social impacts of climate change over time.
@@
To effectively analyze the aforementioned impacts of different decarbonization strategies, it is essential to understand how key variables in our model evolve under each scenario and the rationale behind these changes. This part delves into the specific modifications of variables such as the enterprise low-carbon production efficiency coefficient $\left(k_{t}\right)$, production cost per unit ($c_{t}$), original productivity coefficient ($\left.\beta_{t}\right)$, and carbon price $\left(B_{t}\right)$ for each of the decarbonization scenarios: Immediate, Quick, Slow, and No Decarbonization. By doing so, we can better comprehend the short and long-term effects of the transition on firm’s profitability and identify the optimal transition investment ratio $(\alpha)$ for the firm in each case.
@@
Immediate Decarbonization. In the Immediate Decarbonization scenario, enterprises make substantial investments in low-carbon technologies immediately. This aggressive approach aims at achieving rapid decarbonization and substantial reductions in carbon emissions in the shortest possible time. As a result, the enterprise’s low-carbon production efficiency coefficient $\left(k_{t}\right)$ starts at a very high level due to a substantial initial level of investment, reflecting major improvements in production efficiency of the firm. Over time, as the immediate benefits are realized, $k_{t}$ decreases gradually. The production cost per unit $\left(c_{t}\right)$ of the firm decreases quickly as economies of scale and technological efficiencies are realized from the new low-carbon technologies. The original productivity coefficient $\left(\beta_{t}\right)$ increases initially due to enhanced production capabilities of the firm but stabilizes over time as the effects of initial investments are more fully integrated. The carbon price ($B_{t}$) remains at a very high level due to stricter regulations and market adjustments that favor low-carbon outputs.
@@
Quick Decarbonization. The Quick Decarbonization scenario involves rapid but slightly less aggressive investments compared to the Immediate scenario. Enterprises aim at achieving substantial decarbonization within a relatively short time frame but spread their investments over a few periods of time to balance costs and benefits. Consequently, the enterprise’s low-carbon production efficiency coefficient $\left(k_{t}\right)$ increases quickly but not as sharply as in the case of the Immediate scenario. The efficiency gains are substantial for the firm but spread over a slightly longer period of time. The production cost per unit $\left(c_{t}\right)$ decreases rapidly but not as dramatically, reflecting more moderate cost reductions. The original productivity coefficient $\left(\beta_{t}\right)$ shows a steady increase due to technological improvements and better resource allocation by the firm. The carbon price $\left(B_{t}\right)$ keeps at a relatively high level, driven by increasing regulatory pressures and market responses.
@@
Slow Decarbonization. In the Slow Decarbonization scenario, enterprises adopt a gradual approach to transitioning to low-carbon technologies. This approach spreads investments over a longer period of time, allowing enterprises to manage costs and mitigate risks while steadily reducing carbon emissions. As such, the enterprise’s low-carbon production efficiency coefficient $\left(k_{t}\right)$ increases slowly, reflecting incremental improvements in production efficiency for the firm as investments are made gradually. The production cost per unit $\left(c_{t}\right)$ decreases steadily but slowly, indicating progressive cost reductions over time. The original productivity coefficient $\left(\beta_{t}\right)$ shows a gradual increase, correlating with incremental enhancements in productivity as new technologies are adopted by the firm over time. The carbon price $\left(B_{t}\right)$ is set as a relatively low level, aligning with gradual regulatory changes and market adjustments.
@@
No Decarbonization. The No Decarbonization scenario assumes that enterprises do not make any major investments in low-carbon technologies. Instead, they continue with their current practices, resulting in minimal changes to carbon emissions and production efficiency. Consequently, there is no need to find the optimal transitional investment ratio under this scenario, as the transitional investment ratio, $\alpha$, should always be zero.
@@
By analyzing these different scenarios numerically, we can observe how different levels of investment and urgency in transitioning to low-carbon technologies affect key variables of the framework over time. This analysis helps us understand the trade-offs between immediate costs and long-term benefits, guiding enterprises in making informed decisions about their decarbonization strategies. The goal is to determine the optimal value of $\alpha$ that maximizes the overall profit across multiple periods, ensuring both short-term (3-year observation period) profitability and long-term (6-year observation period) sustainability. We will not consider the scenario of ‘No Decarbonization’ in this part, as there is no need to determine the optimal transitional investment ratio $\alpha$ under this scenario.
@@
In the Quick Transition scenario, enterprises adopt low-carbon technologies at a moderately fast pace. This approach aims at achieving a rapid improvement in production efficiency, though at a slightly slower rate than in the Immediate scenario. As a result, the initial efficiency gains are significant but not as drastic, allowing the enterprise to avoid potential issues related to over-investment. Given the initially low value of $k$ (starting at 1.5), there is an urgent need to enhance production efficiency quickly. To address this, the enterprise makes a transition investment before the second year, resulting in a sharp increase in $k$ to 5 by the second year. This value then gradually declines each subsequent year, reflecting the natural efficiency losses due to wear and tear. Production costs per unit begin at a relatively high level (2.4), creating a strong need for transition investment to reduce costs. With the Quick Decarbonization approach, $c$ drops markedly to 1.6 in the second year. However, it then rises slightly in the following years, likely due to normal depreciation and the ongoing costs associated with maintaining the new technology. Similarly, the original productivity coefficient $\beta$ starts at a lower level (0.6), emphasizing the necessity for improvement. Due to the transition investment, $\beta$ reaches a peak of 0.9 in the second year, reflecting increased productivity. Over time, it gradually decreases, aligning with the adaptation phase and the natural reduction in productivity gains. Carbon pricing $B$ is maintained at a high but stable level (3) in the Quick Decarbonization scenario. This provides a continuous incentive for a firm to adopt low-carbon technologies, without imposing excessive financial burdens on the firm.